OpenStudy (anonymous):

how about this one? (9x3+54x2+81x)−−−−−−−−−−−−−−−√/(x3−4x2+4x

OpenStudy (anonymous):

First take the derivative and set equal to zero: f'(x) = 27x^2 -108x +81 = 0 Divide by 27 to get x^2 - 4x + 3 =0 factor: (x-1)(x-3) = 0 So the critical points are x = 1 and 3. However, x=3 is not in the interval so we reject it. Now this is a continuous function on a closed, bounded interval. So by the Extreme value Thm, it must have both a max and min but they may occur at the endpoints. So see what f(x) is at the endpoints and at the critical point: f(-6) = -4361 f(1) = 49 f(2) = 31 Comparing, we see that the max of f occurs at x = 1 and the min of f occurs at x = -6. [ If you had tested the critical point with the 2nd derivative test you would have gotten: f''(x) = 54x - 108, and at x=1 this is negative. That tells us that x=1 is at least a relative max.] Note: To avoid confusion on these postings, it is best to indicate powers by using the ^ symbol eg: x-cubed is typed x^3.